Solve for x:
20/(x+5)=(x+5)/x

Answer by tommyt3rd(5050)

20/(x+5)=(x+5)/x
(x+5)^2=20x
x^2+10x+25=20x
x^2-10x+25=0
(x-5)^2=0
x=5